OpenStudy (anonymous):

simplify (x^4 over y^2)^2

OpenStudy (anonymous):

(x^8)/(y^4)

OpenStudy (stormfire1):

\[(\frac{x^4}{y^2})^2\]\[=\frac{x^8}{y^4}\]